Completely lost on this question:
x2-3 x
---- = --- (That's x(ex)squared minus 3 over 5 x(ex)plus 2 equals x over 5)
5x+2 5
I think i have to cross distribute for ex. 5(x2-3) and x(5x+2).
Is that right? Then what do i do?!
Please help!
Thanks!

Answer by CharlesG2(834)

Completely lost on this question:
x2-3 x
---- = --- (That's x(ex)squared minus 3 over 5 x(ex)plus 2 equals x over 5)
5x+2 5
I think i have to cross distribute for ex. 5(x2-3) and x(5x+2).
Is that right? Then what do i do?!
Please help!
Thanks!

(x^2 - 3)/(5x + 2) = x/5
cross multiply
(x^2 - 3) * 5 = (5x + 2) * x
distribute 5 on left and x on right
5x^2 - 15 = 5x^2 + 2x
bring like terms together
5x^2 - 5x^2 - 2x = 15
-2x = 15
x = -15/2
check:
x^2 - 3 = 225/4 - 3 = 225/4 - 12/4 = 213/4
5x + 2 = -75/2 + 2 = -75/2 + 4/2 = -71/2 = -142/4
(so that it similar to above line)
(213/4)/(-142/4) = 213/4 * -4/142 = -213/142 = (3 * -71)/(2 * 71) = -3/2
x/5 = (-15/2)/5 = -15/2 * 1/5 = -15/10 = -3/2, yes